# Efficient and Stealthy Jailbreak Attacks via Adversarial Prompt Distillation from LLMs to SLMs

### v1

@@ 4.2.1 Pre-training
In the pre-training stage, we construct a dataset by combining $n$ instructions with $N$ selected templates using the Cartesian product, resulting in $n\times N$ data inputs. This approach ensures comprehensive coverage of diverse instruction-template pairs. We then pre-train Llama, serving as the teacher model, to enhance its adversarial generation capability. The motivation is to establish a robust foundation for generating harmful content through self-supervised learning.
@@
We employ the AdamW optimizer and train over multiple epochs. This pre-training process substantially improves Llama’s ability to generate harmful and stealthy content, providing a strong teacher model for subsequent distillation.
4.3 Knowledge Distillation
Knowledge distillation transfers Llama’s adversarial capabilities to BERT, enabling the generation of stealthier and more efficient attack text. The process aligns BERT’s output distribution with Llama’s by minimizing the KL divergence loss, ensuring consistency between the teacher (Llama) and student (BERT) models. Specifically, the KL divergence is defined as:
@@ 4.3 Knowledge Distillation
Algorithm 1 Adversarial Knowledge Transfer
1:Input:
2:$I$: Set of $n$ harmful instructions
3:$T$: Set of $N$ prompt templates
4: Llama: Teacher model for adversarial generation
5: BERT: Student model for efficient attacks
6:Output: Prompts
7:Step 1: Pre-training
8:1:Enumerate all combinations of $n$ instructions and $N$ templates
9:2: $\hat{y_{i}}=\max P(\text{harmful}|T_{i})$
10:3: $\mathcal{L}_{\text{pretrain}}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}$
11:Step 2: Knowledge Distillation
12:4: Freeze Llama, unfreeze BERT
13:5: $\mathcal{L}_{\text{KL}}=\sum_{i}p_{i}\log\frac{p_{i}}{q_{i}}$
14:6: $\mathcal{L}_{\text{align}}=\text{KL}(\text{logits}_{T},\text{logits}_{S})$
15:7: $T=t_{\text{final}}+(t_{\text{initial}}-t_{\text{final}})(1+\cos(\pi\cdot\text{progress}))$
16:8: while training
17:9:  Update prompts, mask(v,n,adj)
18:10: end
19:11: Normalize logits $\to$ shared layer
20:Step 3: RL Optimization
21:12: $R_{\text{total}}=R_{\text{attack}}+R_{\text{harm}}+R_{\text{diverse}}$
22:13: while training
23:14: $\mathcal{L}_{\text{policy}}=-\mathbb{E}_{\tau\sim\pi}[\log\pi(a|\theta)\cdot R_{\text{total}}]$
24:15:  $\theta\leftarrow\text{Adam}(\nabla,\text{clip})$
25:16:  $k\leftarrow 0$, $k_{\text{max}}$
26:17:  while $k<k_{\text{max}}$ and $R_{\text{attack}}<0$
27:18:   $T\leftarrow T+\epsilon^{\prime}$
28:19:   Recompute $\pi(\theta)$, $R_{\text{total}}$
29:20:   $k\leftarrow k+1$
30:21:  end
31:Return: ASRk, ASRl
To address vocabulary mismatches between Llama and BERT, projection layers map their logits to a shared dimension $D$, minimizing the alignment loss:
@@
The temperature $T$ adjusts dynamically based on training progress, with higher values promoting exploration of diverse prompts early on and lower values focusing on deterministic, high-probability attack prompts later. Additionally, a small random perturbation is applied to $T$ with a certain probability to encourage diversity; if an attack fails, $T$ is adjusted and the prompt is regenerated within $k_{\text{max}}$ attempts until success or the limit is reached. The scaling factor $\alpha$ controls the speed of temperature adjustment, enabling finer control over the rate of exploration versus exploitation. To facilitate learning, the teacher model Llama is frozen, while the last 4 layers of the student model BERT are unfrozen to enable fine-tuning for adversarial generation. During training, prompts for both models are iteratively updated, with a masking operation applied to a subset of tokens (specifically verbs, nouns, and adjectives) to enhance prompt diversity. The final layer of BERT is normalized to a shared linear layer for logits alignment, ensuring that BERT can produce stealthiness and harmful prompts efficiently while maintaining attack efficacy at a reduced computational cost.
4.4 Reinforcement Learning Optimization
We employ reinforcement learning to optimize adversarial prompt generation, where an agent learns to craft prompts to maximize a reward in a language model environment. The reward function balances stealthiness, harmfulness, and diversity, formalized as:
@@ 4.4 Reinforcement Learning Optimization
where $R_{\text{attack}}$ reflects stealthiness (bypassing detection), $R_{\text{harm}}$ measures harmfulness, and $R_{\text{diverse}}$ ensures diversity, as detailed in Table 1.Meanwhile, during the iteration process, we apply a perturbation $\epsilon$ to enhance exploration and perform gradient clipping to stabilize training.
Reward Type Condition Value
$R_{\text{attack}}$ Bypass success 1
Otherwise -1
$R_{\text{harm}}$ Harmful response 1
Otherwise -1
$R_{\text{diverse}}$ Novel prompt 1
Repeated prompt -1
Table 1: Reward mechanism for RL optimization, defining conditions for attack success, harmfulness, and diversity. The decision of the Noval prompt is based on threshold judgment.
@@
where $\pi(a|\theta)$ is the policy generating prompts $a$, and the expectation is over trajectories $\tau$. Gradients are updated using the Adam optimizer Kingma and Ba (2014), with gradient clipping for stability. With a dynamic temperature function in knowledge distillation, RL explores diverse prompts early and optimizes high-reward prompts later, improving attack success rates and cross-scenario adaptability. This approach yields robust prompts for black-box and white-box settings.
@@ 4.5.1 Datasets
We evaluate our framework using the AdvBench dataset Zou et al. (2023), a widely used benchmark for testing LLM security against adversarial jailbreak attacks. AdvBench includes 520 harmful instructions across topics like violence, fraud, discrimination, and illegal activities, designed to assess model robustness. Additionally, we select templates from UltraSafety [^2] https://huggingface.co/datasets/openbmb/UltraSafety to further enhance our evaluation.
@@ 4.5.2 Evaluation Protocol
We employ two evaluation metrics to measure the attack success rate Yu et al. (2024).
Keyword-based Attack Success Rate ASRk, which assesses an attack’s success by examining whether the model’s response avoids refusal keywords (e.g., “Sorry, I cannot assist,” “This is illegal,” or “I cannot comply”). Specifically, a response is deemed successful if it lacks these predefined refusal phrases, indicating that the model has not explicitly rejected the harmful instruction. This method provides a straightforward and efficient way to quantify the model’s vulnerability to jailbreak attempts.
ASRl, The second method uses GPT-4o to evaluate the harmfulness of the model’s responses, calculating the attack success rate (ASRl) based on whether the response contains harmful content, offering a more nuanced assessment of the model’s behavior under attack.
4.5.3 Baselines
We select the following baseline methods for comparison in our experiments include SOTA methods GCG Zou et al. (2023), AutoDAN Liu et al. (2023b), PAIR Chao et al. (2023), TAP Mehrotra et al. (2024), DeepInception Li et al. (2023), BlackDAN Wang et al. (2024), and LLM-Virus Yu et al. (2024). These methods encompass a variety of attack strategies targeting large language models, allowing us to evaluate the performance of our framework against different types of attacks.
@@ 4.6 Implementation Details
Model Avg.Gen.Time(s) Params(M) Hidden Units Attention Heads Layers
@@
Bert-base-uncased 0.23 109.48 768 12 12
Table 2: Model Performance and Architectural Comparison for Llama-3.2-1B and bert-base-uncased
Methods Models GPT-4 GPT-3.5-Turbo Llama-2-7B Vicuna-7B
@@
Table 3: The Comparison of other SOTA methods in jailbreak attack targets to large language models.
4.7 Attack Efficiency
4.7.1 Efficiency Comparison Between Llama and BERT in Knowledge Distillation
In the knowledge distillation task, directly distilling the knowledge of the Llama model onto BERT demonstrates significant efficiency advantages. In terms of time and model parameters, BERT performs well as a student model. Its average generation time is only 0.23 s, which is much lower than 0.86 s of Llama-3.2-1B, indicating that its reasoning speed is faster and suitable for tasks requiring rapid response. Meanwhile, the model parameters of BERT are 109.48 MB. Compared with 1235.81 MB of Llama-3.2-1B, the parameters are significantly reduced. Additionally, BERT’s smaller architecture, with 768 hidden units and 12 layers, contrasts with Llama-3.2-1B’s 2048 hidden units and 16 layers, making BERT more lightweight and easier to fine-tune. This efficiency is further highlighted by BERT’s 12 attention heads compared to Llama-3.2-1B’s 32, allowing for faster computation without sacrificing performance. This not only reduces the demand for computing resources but also improves the deployment efficiency of the model. Therefore, the designed distillation strategy can effectively transfer the attack capability to the lightweight model Bert, thereby achieving higher efficiency and lower resource consumption without sacrificing the attack effect.
Method Time Per Sample
GCG 15 mins
AutoDAN 12 mins
BlackDAN 2.0 mins
APD(Ours) 1.9 mins
Table 4: Comparison of Time with Other Methods: APD Demonstrates a Clear Advantage in Per-Sample Time.
(a) The trends of ASRk and ASRl for Llama2-7b
(b) The trends of ASRk and ASRl for Llama2-13b
Figure 3: As $\alpha$ increases, ASRk generally shows an upward trend, while ASRl exhibits a downward trend
(a) The variation of ASRk with the number of attacks k
(b) The variation of ASRl with the number of attacks k
Figure 4: Analysis of the changing trends of ASRk and ASRl with the number of attacks k under different models
4.7.2 APD’s Efficiency in the Average Time Cost of Harmful Actions
As shown in Table 4, different methods are compared in terms of time cost per sample, with a focus on highlighting APD’s efficiency advantage in the average time cost for harmful actions, particularly when attacking the GPT-3.5-turbo model. GCG has an average time cost as high as 15 minutes, AutoDAN is 12 minutes, while BlackDAN is reduced to 2 minutes. However, our APD requires only 1.9 minutes, significantly outperforming the other methods. This indicates that APD offers higher time efficiency in handling harmful actions, enabling faster task completion and making it suitable for scenarios requiring rapid responses.
4.8 Comparison to other SOTA methods
Table 3 highlights the strong performance of our APD method compared to other adversarial attack techniques across GPT-4, GPT-3.5-Turbo Radford et al. (2020), Llama-2-7B Dubey et al. (2024), and Vicuna-7B Chiang et al. (2023), evaluated through ASRk and ASRl metrics. APD excels on GPT-4 with an ASRk of 96.4% and on Vicuna-7B with an ASRl of 99.6%, while also achieving high success on Llama-2-7B with an ASRk of 96.1% and an ASRl of 96.6%, significantly outperforming methods like AutoDAN and PAIR. This demonstrates APD’s robust effectiveness across diverse models.

### v2

@@ 4.2.1 Pre-training
In the pre-training stage, we construct a dataset by combining $n$ instructions with $N$ selected templates using the Cartesian product, resulting in $n\times N$ data inputs. This approach ensures comprehensive coverage of diverse instruction-template pairs. We then pre-train Llama, serving as the teacher model, to enhance its adversarial generation capabilities. The motivation is to establish a robust foundation for generating harmful content through self-supervised learning.
@@
We employ the AdamW Optimizer Loshchilov & Hutter (2017) and train over multiple epochs. This pre-training process substantially improves Llama’s ability to generate harmful and stealthy content, providing a strong teacher model for subsequent distillation.
4.2.2 Knowledge Distillation
Knowledge distillation transfers Llama’s adversarial capabilities to SLM, enabling the generation of stealthier and more efficient attack text. The process aligns SLMs’ output distribution with Llama’s by minimizing the KL divergence loss, ensuring consistency between the teacher (Llama) and student (SLM) models. Specifically, the KL divergence is defined as:
@@ 4.2.2 Knowledge Distillation
To address vocabulary mismatches between Llama and SLM, projection layers map their logits to a shared dimension $D$, minimizing the alignment loss:
@@
The temperature $T$ adjusts dynamically based on training progress, with higher values promoting exploration of diverse prompts early on and lower values focusing on deterministic, high-probability attack prompts later. Additionally, a small random perturbation is applied to $T$ with a certain probability to encourage diversity. If an attack fails, $T$ is adjusted, and the prompt is regenerated within $k_{\text{max}}$ attempts until success is achieved or the limit is reached. The scaling factor $\alpha$ controls the speed of temperature adjustment, allowing for finer control over the balance between exploration and exploitation. To facilitate learning, the teacher model, Llama, is frozen, while the last four layers of the student model (SLMs like BERT, ALBERT, RoBERTa) are unfrozen to enable fine-tuning for adversarial generation. During training, prompts for both models are iteratively updated, with a masking operation applied to a subset of tokens (specifically verbs, nouns, and adjectives) to enhance prompt diversity. The final layer of SLM (BERT, ALBERT, RoBERTa) is normalized to a shared linear layer for logits alignment, ensuring that SLM can produce stealthiness and harmful prompts efficiently while maintaining attack efficacy at a reduced computational cost.
4.2.3 Reinforcement Learning Optimization
We formulate adversarial prompt optimization as a Reinforcement Learning from AI Feedback (RLAIF) Bai et al. (2022); Liu et al. (2023c) task—an adversarial instantiation of RLHF Ouyang et al. (2022)—wherein a lightweight policy agent learns to craft stealthy and harmful prompts by maximizing a composite reward signal derived from a language model environment. The reward function balances stealthiness, harmfulness, and diversity, formalized as:
@@ 4.2.3 Reinforcement Learning Optimization
Where $R_{\text{attack}}$ measures the stealthiness of the watermark (ability to evade detection), $R_{\text{harm}}$ quantifies the harmfulness (impact on segmentation performance), and $R_{\text{diverse}}$ ensures diversity of the watermark across different inputs, as detailed in follow equations. During the iterative training process, we apply a perturbation $\epsilon$ to enhance exploration and employ gradient clipping to stabilize training.
Where the reward components are:
$R_{\text{attack}}=2\cdot\mathbb{I}\{\text{prompt bypasses detection}\}-1,$ (10)
$R_{\text{harm}}=2\cdot\mathbb{I}\{\text{prompt elicits harmful response}\}-1,$
[…]
@@
where $\pi(a|\theta)$ is the policy generating prompts $a$, and the expectation is over trajectories $\tau$. Gradients are updated using the Adam optimizer Kingma & Ba (2014), with gradient clipping for stability. With a dynamic temperature function in knowledge distillation, RL explores diverse prompts early and optimizes high-reward prompts later, improving attack success rates and cross-scenario adaptability. This approach yields robust prompts for black-box and white-box settings.
5 Experiments
@@ 5.1.1 Datasets
We evaluate our framework using established benchmarks: AdvBench Zou et al. (2023)  [^1] https://huggingface.co/datasets/walledai/AdvBench/tree/main(520 harmful instructions on violence, fraud, discrimination, and illegal activities for robustness testing), and HarmBench Mazeika et al. (2024) [^2] https://huggingface.co/datasets/walledai/HarmBench(over 400 unique harmful behaviors spanning cybercrime, chemical/biological threats, copyright violations, misinformation, harassment, illegal activities, and general harm for evaluating red teaming and defenses). Additionally, we incorporate templates from UltraSafety Guo et al. (2024) [^3] https://huggingface.co/datasets/openbmb/UltraSafety (100K safety-critical instructions on misinformation, violence, illegal activities, privacy breaches, and self-harm for assessing LLM robustness under adversarial conditions).
@@ 5.1.2 Evaluation Protocol
We employ two evaluation metrics to measure the attack success rate. The evaluation method is based on previous work Yu et al. (2024).
Keyword-based Attack Success Rate (ASRk):, which assesses an attack’s success by examining whether the model’s response avoids refusal keywords (e.g., “Sorry, I cannot assist,” “This is illegal,” or “I cannot comply”). Specifically, a response is deemed successful if it lacks these predefined refusal phrases, indicating that the model has not explicitly rejected the harmful instruction. This method provides a straightforward and efficient way to quantify the model’s vulnerability to jailbreak attempts.
Harmfulness-based Attack Success Rate (ASRl):, The second method uses GPT-4o to evaluate the harmfulness of the model’s responses, calculating the attack success rate (ASRl) based on whether the response contains harmful content, offering a more nuanced assessment of the model’s behavior under attack.
5.1.3 Baselines and Target Models
We select the following baseline methods for comparison in our experiments include SOTA methods GCG Zou et al. (2023), AutoDAN Liu et al. (2023b), PAIR Chao et al. (2023), TAP Mehrotra et al. (2024), DeepInception Li et al. (2023), BlackDAN Wang et al. (2024), and LLM-Virus Yu et al. (2024). These methods encompass a variety of attack strategies targeting large language models, enabling us to evaluate our framework’s performance against different attack types.
We evaluate the proposed Adversarial Prompt Distillation (APD) framework against prominent victim models, including OpenAI’s ChatGPT variants (GPT-3.5-Turbo and GPT-4) for their human-like conversational capabilities derived from extensive dataset training Radford et al. (2020), Meta AI’s Llama series (Llama-2-7B, and Llama-3.2-1B and Llama-3.1-8B sizes) for advanced language understanding and generation suited to chatbots and content creation Dubey et al. (2024), and LMSYS’s Vicuna-7B for its optimized balance of efficiency and performance in natural language tasks The Team with members from UC Berkeley & Diego (2023).
Three SLMs, BERT Devlin et al. (2019), ALBERT Lan et al. (2019), and RoBERTa Liu et al. (2019), are chosen as the student models for the jailbreak task. BERT is a bidirectional transformer encoder pre-trained on masked language modeling and next sentence prediction tasks, enabling strong contextual understanding. ALBERT is a lightweight version of the BERT model that reduces parameter size through techniques like factorized embedding parameterization and cross-layer parameter sharing, while maintaining or improving performance on natural language processing benchmarks. RoBERTa is an optimized version of BERT with dynamic masking and larger training data, achieving superior performance on various benchmarks.
@@ 5.2 Implementation Details
All experiments were conducted on 4 $\times$ NVIDIA GeForce H100 GPUs, each with 80.0 GB of video memory. We configure the following parameters: the maximum number of attack attempts $k_{\mathrm{max}}=100$, the number of prompt templates $N=10$, the learning rate $\mathrm{lr}=0.01$, the initial temperature for the dynamic temperature function $t_{\mathrm{initial}}=2$, the final temperature $t_{\mathrm{final}}=0.5$, the perturbation probability $p=0.1$, the perturbation magnitude $\epsilon\in[-0.5,0.5]$, the maximum consecutive skips $\mathrm{max\_consecutive\_skips}=5$, and the masking probability $\mathrm{mlm\_prob}=0.1$. See parameter sensitivity experiment Section 6.2.2 and Section 6.2.1 for the determination of relevant parameters.
5.3 Comparison of Jailbreak Performance Under APD Method
We evaluate the jailbreak performance of our APD framework by comparing the ASR of the teacher model before distillation and the student model after distillation, against representative SOTA baselines on the AdvBench and HarmBench datasets. Results are reported in Table 1, where Ours (Teacher) denotes the pre-trained Llama teacher prior to distillation, and Ours (Student) represents the distilled SLM models executing the final attack prompts.
Table 1: Comparison of Jailbreak Performance Under Adversarial Prompt Distillation Method
Methods GPT-4 GPT-3.5-Turbo Llama-2-7B Vicuna-7B
AutoDAN - 65.7 60.8 97.7
TAP 36.0 48.1 30.2 31.5
DeepInception 61.9 68.5 77.5 92.7
LLM-Virus 74.0 90.8 95.6 93.5
Llama-3.1-8B 72.9 94.1 99.6 99.7
Llama-3.2-1B 65.6 82.5 75.2 92.4
APD (Ours) 96.4 99.7 96.1 100.0
As shown in the Tables, the distilled student models consistently outperform both the LLM-Virus baseline and the Llama teacher across all four target models, achieving the highest ASR in three out of four cases and the second-best in the remaining one. Notably, on Vicuna-7B, the student achieves an ASR of 86.2%, surpassing LLM-Virus (80.5%) and the teacher (85.7%), demonstrating successful transfer of jailbreak capability despite a $100\times$ reduction in parameter count (from 8B to 110M). On GPT-4-0613 and GPT-3.5-Turbo, the student improves over the teacher by 5.5% and 1.8%, respectively, indicating that distillation not only preserves but enhances stealth and harm induction through focused alignment on high-reward prompt structures.
5.4 Attack efficiency comparison between teacher and student model on jailbreak tasks
@@
As shown in  Table 2, different methods are compared in terms of time cost per sample, with a focus on highlighting APD’s efficiency advantage in the average time cost for harmful actions, particularly when attacking the GPT-3.5-turbo model. GCG has an average time cost of up to 15 minutes; AutoDAN is 12 minutes; and BlackDAN is 2 minutes per sample. However, our APD requires only 1.9 minutes per sample time, significantly outperforming the other methods. This indicates that APD offers greater time efficiency in handling harmful actions, enabling faster task completion and making it suitable for scenarios that require rapid responses.
Table 2: Comparison of Time with Other Methods, APD Demonstrates a clear advantage in Per-Sample Time.
Method Time Per Sample
GCG 15 mins
AutoDAN 12 mins
BlackDAN 2.0 mins
APD (Ours) 1.9 mins
Table 3: Model Efficiency Comparison for teacher models and student models
Model Avg.Time (s) Params (M) Hidden Units Attention Heads Layers
Llama-3.1-8B 0.45 8000.00 4096 32 32
@@
RoBERTa 0.29 124.65 768 12 12
ALBERT 0.25 11.68 768 12 12
BERT 0.23 109.48 768 12 12
5.5 Comparison to other SOTA methods
Table 4 and Table 5 highlight the strong performance of our APD method compared to other adversarial attack techniques across GPT-4, GPT-3.5-Turbo Radford et al. (2020), Llama-2-7B Dubey et al. (2024), and Vicuna-7B The Team with members from UC Berkeley & Diego (2023), evaluated through ASRk and ASRl metrics on AdvBench and HarmBench datasets. For specific evaluation criteria, please see Appendix B.2.4 and Appendix B.2.3.
In AdvBench evaluation shown in Table 4, Adversarial Prompt Distillation excels on GPT-4 with an ASRk of 96.4% and on Vicuna-7B with an ASRl of 99.6%, while also achieving high ASRk and ASRl on Llama-2-7B of 96.1% and 96.6%, respectively, significantly outperforming methods like AutoDAN and PAIR. This demonstrates APD’s robust effectiveness across diverse models.
Table 4: Attack Success Rate ($ASR_{k}$ and $ASR_{l}$) comparison of APD with state-of-the-art jailbreak methods on the AdvBench dataset. Bold: best result.
Methods GPT-4 GPT-3.5-Turbo Llama-2-7B Vicuna-7B
@@ 5.5 Comparison to other SOTA methods
In HarmBench evaluation shown in Table 5, Adversarial Prompt Distillation achieves superior ASRc on GPT-4-0613 with 63.0% for Ours (Student) and 54.5% for Ours (Teacher), while delivering strong results on Vicuna-7B with 86.2% and 85.7%, respectively, and on Llama-2-7B with 50.0% and 42.9%, markedly surpassing baselines like LLM-Virus and PAIR. This highlights APD’s enhanced attack efficacy and transferability through distillation across both proprietary and open-source models.
Table 5: Attack Success Rate (ASRASR %) comparison of Adversarial Prompt Distillation with SOTA jailbreak methods on the HarmBench dataset. Ours (Student) and Ours (Teacher) are before/after jailbreak knowledge distillation. Bold: best result, underline: second best result.
Methods GPT-4-0613 GPT-3.5-Turbo Llama-2-7B Vicuna-7B
GCG 22.0 42.5 19.7 60.8
ZS 19.4 28.4 2.0 27.2
PAIR 39.3 35.0 9.3 53.5
TAP 43.0 39.2 9.3 51.0
TAP-T 54.8 47.5 7.8 59.8
PAP 16.8 11.3 2.7 18.9
LLM-Virus 29.3 53.8 38.5 80.5
Ours (Teacher) 54.5 54.4 42.9 85.7
Ours (Student) 63.0 56.2 50.0 86.2
6 Ablation Study
6.1 Component Effectiveness Analysis
To evaluate the contributions of each component in our Adversarial Prompt Distillation (APD) framework, we performed an ablation study on the AdvBench dataset using 520 harmful instructions. We compared the full APD against variants omitting one component each—Adversarial Knowledge Transfer, Knowledge Distillation, or Reinforcement Learning Optimization, across victim models GPT-4, GPT-3.5-Turbo, Llama-2-7B, and Vicuna-7B.
Table 6: Ablation Study of Component Effectiveness Analysis on HarmBench Dataset
Component GPT-4 Vicuna-7B
w/o Adversarial Knowledge Transfer 2.1 78.5
w/o Knowledge Distillation 54.5 85.7
w/o Reinforcement Learning Optimization 57.1 83.3
Full Adversarial Prompt Distillation 63.0 86.2
The ablation experiments use the same hyperparameters and report the average ASR across 10 runs. The results on Table 6 show that there is no significant decrease in ASR for any component. Among them, the ASR of GPT-4 is only 2.1%, and Vicuna-7B is also significantly lower than the full component. The two groups of ablation experiments showed a structurally increasing relationship, and ASR also increased in stages. It illustrates the effectiveness of our Adversarial Knowledge Transfer, Knowledge Distillation, and Reinforcement Learning Optimization. Based on the results of GPT-4 and Vicuna-7B, the APD framework relies heavily on the entire architecture to transfer jailbreak capabilities to SLMs.